OpenStudy (anonymous):

Is the line through points P(-8, 4) and Q(9, -6) parallel to the line through to the line through points R(5, -3) and S(0, 2)? A) Yes, the lines have equal slopes. B) No, one line has slope, while the other has no slope. C) Yes, the lines are both vertical. D) No, the lines have unequal slopes.

OpenStudy (mertsj):

Can you find the slopes?

OpenStudy (anonymous):

Hello my son.

OpenStudy (anonymous):

I shalt explain you how to find the slope.

OpenStudy (anonymous):

@Ruhhr thou there?

OpenStudy (mertsj):

\[\frac{-6-4}{9-(-8)}=\frac{-10}{17}\]

OpenStudy (anonymous):

You art correct, Merts.

OpenStudy (mertsj):

\[\frac{-2-(-3)}{0-5}=\frac{1}{-5}\]

OpenStudy (mertsj):

Well, I see the slopes are not the same. So....I'll leave the conclusion to you.